The length of a rectangle is 3n+2 and its width is n-1. The perimeter of the rectangle is twice the sum of its length and its width. Write an expression that represents the perimeter.
Answer by mananth(16946)

Perimeter = 2(L+W)
Perimeter = 2(3n+2 + n-1)
Perimeter = 2(4n+1)
Perimeter = 2(4n+1) ===> 8n+2